OpenStudy (anonymous):

evaluate 4(x+3)+7= -2(3x+1)

OpenStudy (anonymous):

The first step in this problem is distributing all of the numbers that are being multiplied: 4∗x+4∗3+7=−2∗3x+−2∗1 After this, combine all of the 'x' terms and all of the number terms. The last step is to isolate all of the 'x's onto one side and solve for x!

OpenStudy (anonymous):

4(x+3) +7 = -2 ( 3x+1) First, distribute everything outside the parentheses 4x+12 + 7 = -6x -2 Now combine like terms 12 + 7 = 19 19 + 4x = -6x - 2 Now take all x terms to one side, isolate them 10 x = -21 Now divide by 10 x = -21/10

OpenStudy (anonymous):

Thanks for your help